OpenStudy (anonymous):

help pwease?! confused... Which ordered pair is a solution for this system of inequalities? 2x + 3y > 1 and y < 5/4 x - 1 (3, 9) (-4, 2) (8, -1) (-2, -5)

OpenStudy (anonymous):

just plug in the numbers and so which works out

OpenStudy (helder_edwin):

did u do what @Math_Is_Confusing123 suggested?

OpenStudy (anonymous):

yes but still sort of confused

OpenStudy (anonymous):

sorry im a little slow T.T lmao

OpenStudy (helder_edwin):

for instance (3,9)=(x,y) then the firrst inequality would say \[ 2x+3y=2\cdot3+3\cdot9=6+27=36 \] is 36>1? and the seconde inequality could be rewritten as \[ 0<5/4x-y-1 \] so \[ 5/4x-y-1=5/4\cdot3-9-1=15/4-10=-25/4 \] is -25/4>0?

OpenStudy (anonymous):

OH i think i understand. thanks i just needed some sort of explination ! thank you very much ^^~

OpenStudy (helder_edwin):

u r welcome